# Ctrl-V: Higher Fidelity Video Generation with Bounding-Box Controlled Object Motion

### v2

@@ 3.4 Ctrl-V: Box2Video
Given an initial frame and its corresponding sequence of bounding boxes, our aim is to produce a video clip in which the object positions align with the bounding-box sequence. To achieve this, we introduced Box2Video network, where we adapt ControlNet for the SVD framework. Unlike previous works such as Boximator and TrackDiffusion[5, 35], this motion-control model is trained in a single-stage without requiring additional optimization criteria.
Our Box2Video comprises two main components: an SVD and a ControlNet. The input to the SVD model remains consistent with Equation 2. Regarding the ControlNet, we utilize the image encoder to encoder the bounding-box frames $\mathcal{E}({\bm{f}}_{\text{bbox}})={\bm{b}}\in\mathbb{R}^{N,C^{\prime},H^{\prime},W^{\prime}}$. Additionally, we incorporate an adapter layer to preprocess the bounding-box latents before adding them to the frame latents. The inputs to the ControlNet are ${\mathbb{C}}_{\xi}\big{(}{\bm{b}},\hat{{\bm{z}}}_{t},{\bm{z}}_{\text{pad}}^{(0)},{\bm{c}}^{(0)},t\big{)}$.
The architecture of ControlNet closely resembles the contraction and mid-blocks of the SVD’s U-Net architecture, a concept introduced by [1]. Additionally, at the start of the training phase, ControlNet’s weights are initialized to match those of their mirrored layers. Similar to U-Net, the CLIP encoded token of the initial frame (${\bm{c}}^{(0)}$) and the time-step embedding are also injected into every block within the ControlNet. Finally, the output from each ControlNet block is added to the residual path within the SVD model after zero-convolution computation.
During training, the weights of the SVD model ($\theta$) are frozen, while only the weights in the ControlNet ($\xi$) are updated.
@@ 4 Experimental Analysis and Ablation Studies
All experiments are trained on the training sets and evaluated on the testing sets. To evaluate the quality of the generated videos, we randomly select 200 frames from the testing set of each dataset as the initial frames and generated videos based on them. For each dataset, the results presented in this section are based on analyses of the 200 generations.
Furthermore, each bounding-box predictor discussed in the following section requires conditioning on ONE initial frame, ONE or THREE initial bounding-box frames and ONE final bounding-box/trajectory frame.
@@ 4.1 Datasets
We evaluate the performance of our models across three autonomous-vehicle datasets: KITTI [6], Virtual KITTI 2 (vKITTI) [7], and the Berkeley Driving Dataset (BDD) [8] with Multi-object Tracking labels (MOT2020).
KITTI comprises 22 real-world driving clips with 3D object labelling. vKITTI consists of 5 virtual simulated driving scenes, each offering 6 weather variants, all including 3D object labelling. BDD is a large-scale real-world driving dataset, featuring 1603 2D-labeled sequences of driving clips. Further details on dataset configurations are provided in Appendix A.3.
4.2 Generation Quality
[…]
KITTI Stable Video Diffusion Baseline [11] None 552.7 0.3504 0.4030 13.01
BBox Predictor + Box2Video 1-to-1 467.7 0.3416 0.3241 13.21
BBox Predictor + Box2Video 3-to-1 422.2 0.2981 0.4277 13.85
Teacher-forced Box2Video All 435.6 0.2963 0.4394 14.10
vKITTI Stable Video Diffusion Baseline [11] None 331.0 0.2852 0.5540 16.60
BBox Predictor + Box2Video 1-to-1 400.2 0.3179 0.4714 15.78
BBox Predictor + Box2Video 3-to-1 341.4 0.2645 0.5841 17.60
Teacher-forced Box2Video All 313.3 0.2372 0.6203 18.41
BDD Stable Video Diffusion Baseline [11] None 409.0 0.3454 0.5379 16.99
BBox Predictor + Box2Video 1-to-1 412.8 0.2967 0.5470 17.52
BBox Predictor + Box2Video 3-to-1 373.1 0.3071 0.5407 17.37
[…]
Teacher-forced Box2Video All 348.9 0.2926 0.5836 18.39
Table 1: Generated videos consist of 25 frames at a resolution of $312\times 520$, while reported metrics from this table are evaluated at a resolution of $256\times 410$. The “# Cond. BBox” column reports the number of ground-truth input bounding-box frames used by the generation pipelines. “None” indicates that no ground-truth frames are used, while “All” indicates that all ground-truth bounding-box frames are utilized. If “# Cond. BBox” is $n\text{-to-}m$, then it represents the number of initial bounding-box frames used by the pipeline is $n$ and the number of final bounding-box frames used by the pipeline is $m$. If the reported number of bounding-box frames has a superscript “Traj”, it indicates that those bounding-box frames have been replaced by trajectory frames.
To assess the quality of video generation, we compare videos generated through 3 distinct pipelines: fine-tuned SVD baselines (from initial frame to video), teacher-forced Box2Video generation (from initial frame$+$all bounding-box frames to video), and bounding-box prediction with Box2Video (from initial frame$+$initial bounding-box frames to video).
We evaluate our generation across four metrics: Fréchet Video Distance (FVD) [36], Learned Perceptual Image Patch Similarity (LPIPS) [37], Structural Similarity Index Measure (SSIM) [38] and Peak Signal-to-Noise Ratio (PSNR). These metrics either measure the consistency of frame pixels with the ground truth or the consistency of the frame latents extracted by another network. FVD is an exception; it evaluates the generation distribution against the ground truth’s distribution. It is important to note that while many papers report their best-out-of-K results on these metrics, due to computational constraints, we evaluate our model on a single sample for each input.
The evaluated results are reported in Table 1 and visualizations are available in Appendix C.1. These results indicate that the generation quality improves as we condition on more ground-truth bounding-box frames. Details regarding the metrics and their limitations are discussed in Appendix B.1.
4.3 BBox Predictor: Quantitative Evaluation
@@
KITTI 1-to-1 .$629\pm.212$$763\pm.188$$758\pm.176$$986\pm.012$$994\pm.008$$992\pm.009$
3-to-1 .$795\pm.112$$884\pm.078$$881\pm.082$$986\pm.010$$992\pm.007$$994\pm.005$
vKITTI 1-to-1 .$710\pm.205$$809\pm.171$$828\pm.178$$943\pm.048$$946\pm.046$$997\pm.006$
3-to-1 .$767\pm.131$$853\pm.078$$881\pm.126$$944\pm.039$$948\pm.036$$996\pm.006$
BDD 1-to-1 .$587\pm.214$$712\pm.194$$747\pm.187$$954\pm.047$$955\pm.047$$9$002
@@
3-to-1 .$647\pm.176$$783\pm.156$$784\pm.150$$955\pm.043$$955\pm.042$$997\pm.001$
Table 2: All experiments reported in this table are conditioned on 3 initial bounding-box frames and 1 final bounding-box or trajectory frame. The first three columns show evaluations on the entire generated bounding-box sequence, measuring the alignment scores between our generated bounding box predictions and ground-truth labels. The last three columns focus on testing the auto-encoding capability of the network, evaluating only the first and last frames of the generated sequence. If a dataset’s name is labelled with “Traj” superscript, it indicates that in the experiment reported in that row, the last bounding-box frame is replaced with its trajectory frame.
To evaluate the quality of our bounding-box predictions, we create mask images for both the ground-truth and predicted bounding-box sequences. The mask images are generated by converting the bounding-box frames into binary masks (details can be found in Appendix B.2). We then calculate the generated averaged mask Intersection over Union (maskIoU) scores, averaged mask Precision (maskP) scores, and averaged mask Recall (maskR) scores against the ground-truth bounding-box masks. To assess our predictions, we applied the “best-out-of-K” method, selecting the model with the highest maskIoU score for evaluation. In this instance, K equals 5. We present our findings in Table 2, and demonstrate examples of our bounding-box predictions on each dataset in Appendix C.
In the bounding-box prediction figures, our predictor model achieves the closest alignment with the ground-truth in the first and last frames. This near-perfect alignment is primarily attributed to conditioning the model on the bounding boxes of these key frames. When considering all predicted frames, the alignment scores decrease, as shown by the plotted demonstrations and metric results in Table 2. This is because objects in frames do not move deterministically. The role of the bounding-box predictor is to generate a plausible trajectory for moving objects from the initial bounding box frame to the last.
Despite the disparity between the ground-truth trajectory and the predicted trajectory, our Box2Video consistently generates high-fidelity videos based on either trajectory provided. Further analysis of this aspect is provided in the subsequent sections.
@@ 4.4 Box2Video: Motion Control Evaluation
Figure 3: Illustrations of the generations conditioned on ground truth 3D bounding boxes (2D for BDD) across various datasets. The 2D outlines of bounding boxes are overlayed on top.
Our Box2Video is trained to control object motions through bounding boxes using a teacher-forcing approach. This means that only ground-truth bounding-box frames are provided during the training phase. In this section, we analyze the fidelity of our Box2Video generations to the ground-truth bounding-box conditions. To access the consistency of objects’ locations between our generated content and ground-truth, we compute the average precision of the bounding boxes in the generated frames and the ground-truth frames.
Average precision (AP) scores gauge the alignment of predicted/generated bounding boxes with the ground-truth labeling. In all related prior studies, average precision (AP) scores have been consistently reported. However, it is important to acknowledge that AP scores can vary across studies, depending on the specifics of the task setup. Boximator [5]’s motion control model predicts object locations in the scene, focusing solely on objects with consistent appearances across all frames. Their AP implementation disregards the object locations in the intermediate frames, comparing the objects’ locations only in the final frame. In contrast, TrackDiffusion [35] uses TrackAP for evaluation, employing a QDTrack model [28] to track instances in generated videos and comparing them to ground-truth labels. However, these evaluated datasets have limited instances, and TrackAP requires consistent tracking across frames, making it unsuitable for our project without modifications. Therefore, our AP score differs slightly from those in previous works.
First, we utilize the state-of-the-art object detection tool, YOLOv8 [39], to obtain the objects’ trackings from the generated and ground-truth scenes. Detailed information about the tool and our configurations is reported in Appendix B.3. Next, we match objects in each generated-vs-ground-truth frame pair based on spatial similarity – calculating the intersection over union (IoU) score to determine the similarity in location between objects’ bounding boxes. Our metric disregards object type and tracking IDs equivalence – assuming that objects close in location should naturally have the same type and IDs. Finally, we compute the average precision score following MS COCO protocol [40]. Details are provided in Appendix B.4 and results are listed in Table 4.4. These results indicate that our Box2Video model is particularly adept at adhering to the specified conditions, especially when evaluated with a more lenient metric (i.e., a lower IoU threshold for the AP computation).
@@
Boximator [5] [^1] Boximator only used the last generated frame’s bounding boxes to compute average percision scores. MSR-VTT[41] Web videos 16 0.365 0.521 0.384 -
ActivityNet [42] Human-action 16 0.394 0.607 0.409 -
UCF-101 [43] Human-action 16 0.212 0.343 0.205 -
@@
UCF-101 [43] Human-action 16 0.205 0.326 - -
Table 3: Average Precision scores of our model. While prior works [5, 35] utilize varying evaluation strategies and do not report results on KITTI, vKITTI and BDD, we draw upon their reported performances on alternative datasets to provide a comparative context. Longer videos are associated with decreased quality and lower detection rates, posing an additional challenge for our model.

### v3

@@ 3.4 Ctrl-V: Box2Video
Box2Video is shown on the right in Figure 2 and it aims to generate high-fidelity videos controlled by bounding box frames, such as those generated by the BBox Generator network. Box2Video consists of an SVD backbone for video generation, and an adapted ControlNet module to process the bounding box control signal. ControlNet is a widely used network for controlling image generation. In this work, we modify ControlNet and adapt it to the video diffusion framework (as shown on the right in Figure 2). This architecture allows us to train Box2Video in a single stage without the need for additional optimization criteria, in contrast to previous work such as Boximator and TrackDiffusion (Wang et al., 2024; Li et al., 2024), which require multi-stage learning with extra criteria to train their models.
The SVD component takes two inputs: ${\bm{z}}^{(0)}$ and $\hat{{\bm{z}}}_{t}$. Here, ${\bm{z}}^{(0)}$ is the encoded initial video frame and $\hat{{\bm{z}}}_{t}$ is the encoded full video with $t$ levels of noise added to it. As shown in Figure 2, we process these inputs by padding ${\bm{z}}^{(0)}$ by repeating it along the first dimension before concatenating it with $\hat{{\bm{z}}}_{t}$ to create the final input to the UNet adapter of the SVD. The same input is also sent to the ControlNet module through its own UNet adapter layers. Additionally, ControlNet also receives the encoded bounding box frames, ${\bm{b}}$, as input, through ControlNet adapter layers. Both of these transformed input is then added together before processed by the ControlNet module. The output signal of the ControlNet module then goes through a zero-convolution before being sent to the SVD UNet decoder layers through residual paths as control signal. During training, the weights of the SVD model ($\theta$) are frozen, while only the weights in the ControlNet ($\xi$) are updated.
@@ 4 Experimental Analysis and Ablation Studies
Frame 1 Frame 7 Frame 13 Frame 19 Frame 25
BBox Generator’s
Output: Generated
BBox Trajectory
Box2Video’s
Output:
Generated Frames
Ground Truth
Frames
SVD Baseline
Generated Frames
Figure 3: The first two rows illustrate video samples generated using the Ctrl-V pipeline, with one initial frame, three initial bounding box frames, and one final bounding box frame as input. The first row shows bounding box trajectories from the BBox-generator in pixel space (solid rectangles for predictions, wireframe rectangles for ground truth). The second row presents frames generated by the Box2Video model, conditioned on the BBox-generator’s output. The third row displays ground-truth frames, while the fourth row shows frames generated by the Stable Video Diffusion (SVD) baseline. In the Ctrl-V video, the car with the bright-green bounding box, which initially pokes out its nose in the lane to the left of the ego car, stays beside the ego car in the final frame. Meanwhile, the silver car with the olive bounding box, which starts in the lane to the right of the ego car, speeds off and is replaced by a new car (purple bounding box) entering the frame. These generated frames closely match the car positions seen in the conditioned inputs. In contrast, the SVD-generated video shows the black car on the left accelerating and moving ahead of the ego car, while the silver car remains in the same relative position to the ego car throughout.
For quantitative evaluation, we assess the model’s performance across four driving datasets on three key aspects:
  1. 1.
    The overall visual quality of the generated results (Section  4.3)
  2. 2.
    The alignment of the predicted bounding box trajectories with the ground truth (Section  4.2)
  3. 3.
    The fidelity of the generated objects in the video to the bounding box control signal (Section  4.4)
For visual assessment, Figure 3 and Appendix E showcase sample demonstrations generated by our model. To assess video quality, we randomly select 200 initial frames from each dataset’s testing set and generate videos. The results in this section are based on analyses of these 200 generated videos per dataset. Furthermore, we explored different bounding box conditioning options: one or three initial bounding box frames, followed by a single final bounding box. Additional variations are discussed in Appendix E.8.
@@ 4.1 Datasets
We evaluate the performance of our models across four autonomous-vehicle datasets: KITTI (Geiger et al., 2013), Virtual KITTI 2 (vKITTI) (Cabon et al., 2020), Berkeley Driving Dataset (BDD) (Yu et al., 2020) with Multi-object Tracking labels (MOT2020), and the nuScenes Dataset (Caesar et al., 2019). KITTI comprises 22 real-world driving clips with 3D object labelling. vKITTI consists of 5 virtual simulated driving scenes, each offering 6 weather variants, all including 3D object labelling. BDD is a large-scale real-world driving dataset, featuring 1603 2D-labeled sequences of driving videos. The nuScenes dataset is a large-scale driving dataset that includes 1000 scenes 20-second scenes annotated with 3D bounding boxes, multiple sensor data (lidar, radar and cameras) and map information. Further details on dataset configurations are provided in Appendix C.3.
4.2 BBox Generator: Quantitative Evaluation
Frame 1 Frame 7 Frame 13 Frame 19 Frame 25
Trajectory
Generation #1:
Trajectory
Generation #2:
Figure 4: This figure visualizes two samples of bounding box trajectories generated by the BBox Generator, conditioned on the same set of three initial bounding box frames and one final bounding box frame (solid rectangles represent predictions, and wireframe rectangles represent ground truth). Although the intermediate frames show notable differences, the initial and final frames align closely with the ground-truth bounding boxes.
Method
@@ 4.2 BBox Generator: Quantitative Evaluation
KITTI BBox Generator (ours) 1-to-1 .$\textbf{.629}\pm.212$$\textbf{.758}\pm.176$$\textbf{.763}\pm.188$$\textbf{.986}\pm.012$$\textbf{.994}\pm.008$$\textbf{.992}\pm.009$
Trajeglish-Style .$447\pm.154$$568\pm.172$$679\pm.177$$561\pm.151$$663\pm.150$$789\pm.165$
BBox Generator (ours) 3-to-1 .$\textbf{.795}\pm.112$$\textbf{.881}\pm.082$$\textbf{.884}\pm.078$$\textbf{.986}\pm.010$$\textbf{.992}\pm.007$$\textbf{.994}\pm.005$
Trajeglish-Style .$491\pm.164$$622\pm.173$$691\pm.175$$576\pm.154$$684\pm.149$$784\pm.163$
vKITTI BBox Generator (ours) 1-to-1 .$\textbf{.710}\pm.205$$\textbf{.828}\pm.178$$\textbf{.809}\pm.171$$\textbf{.943}\pm.048$$\textbf{.946}\pm.046$$\textbf{.997}\pm.006$
Trajeglish-Style .$471\pm.171$$578\pm.200$$700\pm.187$$557\pm.171$$628\pm.194$$835\pm.135$
BBox Generator (ours) 3-to-1 .$\textbf{.767}\pm.131$$\textbf{.881}\pm.126$$\textbf{.853}\pm.078$$\textbf{.944}\pm.039$$\textbf{.948}\pm.036$$\textbf{.996}\pm.006$
Trajeglish-Style .$520\pm.162$$630\pm.186$$741\pm.176$$575\pm.154$$657\pm.182$$836\pm.143$
BDD BBox Generator (ours) 1-to-1 .$\textbf{.587}\pm.214$$\textbf{.747}\pm.187$$\textbf{.712}\pm.194$$\textbf{.954}\pm.047$$\textbf{.955}\pm.047$$9$002
Trajeglish-Style .$305\pm.183$$372\pm.213$$658\pm.207$$432\pm.171$$483\pm.192$$840\pm.166$
BBox Generator (ours) 3-to-1 .$\textbf{.647}\pm.176$$\textbf{.784}\pm.150$$\textbf{.783}\pm.156$$\textbf{.955}\pm.043$$\textbf{.955}\pm.042$$\textbf{.997}\pm.001$
Trajeglish-Style .$373\pm.185$$454\pm.206$$686\pm.193$$492\pm.190$$553\pm.208$$842\pm.154$
nuScenes BBox Generator (ours) 1-to-1 .$364\pm.242$$433\pm.278$$\textbf{.740}\pm.186$$\textbf{.983}\pm.013$$\textbf{.985}\pm.0112$$\textbf{.997}\pm.003$
Trajeglish-Style .$\textbf{.405}\pm.202$$\textbf{.506}\pm.220$$661\pm.216$$511\pm.168$$603\pm.172$$789\pm.195$
BBox Generator (ours) 3-to-1 .$\textbf{.827}\pm.150$$\textbf{.892}\pm.120$$\textbf{.906}\pm.099$$\textbf{.983}\pm.013$$\textbf{.985}\pm.012$$\textbf{.998}\pm.003$
Trajeglish-Style .$448\pm.194$$554\pm.213$$695\pm.196$$529\pm.172$$623\pm.177$$791\pm.192$
Table 1: Comparing real and generated bounding boxes. We condition on 1 or 3 initial bounding box frame(s) and 1 final bounding box or trajectory frame. The first three columns show evaluations on the entire generated bounding box sequence, measuring the alignment scores between our generated bounding box generations and ground-truth labels. The last three columns focus on testing the auto-encoding capability of the network, evaluating only the first and last frames of the generated sequence. “BBox Generator” is our method and “Trajeglish-Style” is a baseline inspired from Philion et al. (2023) (see Appendix F for implementation details on this baseline).
Figure 4 showcases two bounding box trajectory samples generated by the  BBox Generator, conditioned on the same initial and final bounding box frames. To evaluate the quality of our bounding box generations, we create mask images for both the ground-truth and generated bounding box sequences. The mask images are generated by converting the bounding box frames into binary masks (details can be found in Appendix D.2). We then calculate the generated averaged mask Intersection over Union (maskIoU) scores, averaged mask Precision (maskP) scores, and averaged mask Recall (maskR) scores against the ground-truth bounding box masks. To assess our bounding box trajectories, we applied the “best-out-of-K” method, selecting the model with the highest maskIoU score for evaluation. In this instance, K equals 5. We compare our results with a baseline referred to as the “Trajeglish-Style” model, an autoregressive GPT-like encoder-decoder that models the bounding box trajectories as a sequence of discrete motion tokens. This baseline is inspired by the work of Philion et al. (2023) with implementation details provided in Appendix F. We present our findings in Table 1, and demonstrate examples of our bounding box generations on each dataset in Appendix E.
In the bounding box generation figures, our generator model achieves the closest alignment with the ground-truth in the first and last frames. This near-perfect alignment is primarily attributed to conditioning the model on the bounding boxes of these key frames. When considering all generated frames, the alignment scores decrease, as shown by the plotted demonstrations and metric results in Table 1. This is because objects in frames do not move deterministically. The role of the bounding box generator is to generate a plausible trajectory for moving objects from the initial bounding box frame to the last.
Despite the disparity between the ground-truth trajectory and the generated trajectory, our Box2Video consistently generates high-fidelity videos based on either trajectory provided. Further analysis of this aspect is provided in the subsequent sections.
4.3 Ctrl-V: Generation Quality
@@
KITTI Stable Video Diffusion Baseline (Blattmann et al., 2023a) None 1118.4 0.4575 0.2919 10.63
Stable Video Diffusion Fine-tuned (Blattmann et al., 2023a) None 552.7 0.3504 0.4030 13.01
Ctrl-V: BBox Generator + Box2Video(Ours) 1-to-1 467.7 0.3416 0.3241 13.21
Ctrl-V: BBox Generator + Box2Video(Ours) 3-to-1 422.2 0.2981 0.4277 13.85
Ctrl-V: Teacher-forced Box2Video(Ours) All 435.6 0.2963 0.4394 14.10
vKITTI Stable Video Diffusion Baseline (Blattmann et al., 2023a) None 922.7 0.3636 0.4740 14.61
Stable Video Diffusion Fine-tuned (Blattmann et al., 2023a) None 331.0 0.2852 0.5540 16.60
Ctrl-V: BBox Generator + Box2Video(Ours) 1-to-1 400.2 0.3179 0.4714 15.78
Ctrl-V: BBox Generator + Box2Video(Ours) 3-to-1 341.4 0.2645 0.5841 17.60
Ctrl-V: Teacher-forced Box2Video(Ours) All 313.3 0.2372 0.6203 18.41
BDD Stable Video Diffusion Baseline (Blattmann et al., 2023a) None 933.6 0.4880 0.3349 12.70
Stable Video Diffusion Fine-tuned (Blattmann et al., 2023a) None 409.0 0.3454 0.5379 16.99
Ctrl-V: BBox Generator + Box2Video(Ours) 1-to-1 412.8 0.2967 0.5470 17.52
Ctrl-V: BBox Generator + Box2Video(Ours) 3-to-1 373.1 0.3071 0.5407 17.37
Ctrl-V: Teacher-forced Box2Video(Ours) All 348.9 0.2926 0.5836 18.39
nuScenes Single-View Stable Video Diffusion Baseline (Blattmann et al., 2023a) None 1179.4 0.5004 0.2877 13.31
Stable Video Diffusion Fine-tuned (Blattmann et al., 2023a) None 316.6 0.2730 0.4787 18.58
Ctrl-V: BBox Generator + Box2Video(Ours) 1-to-1 285.3 0.2647 0.5050 18.93
Ctrl-V: BBox Generator + Box2Video(Ours) 3-to-1 235.0 0.2235 0.5500 20.33
Ctrl-V: Teacher-forced Box2Video(Ours) All 235.5 0.2104 0.5705 23.36
DriveGAN (Kim et al., 2021) None 390.8 - - -
DriveDreamer (Wang et al., 2023b) All 340.8 - - -
Multi-view WoVoGen (Lu et al., 2023) All 417.7 - - -
Drivingdiffusion (Li et al., 2023) All 332.0 - - -
Drive-WM (Lu et al., 2023) None 212.5 - - -
BEVWorld (Zhang et al., 2024) None 154.0 - - -
Panacea (Wen et al., 2024) All 139.0 - - -
Drive-WM (Lu et al., 2023) All 122.7 - - -
DriveDreamer-2 (Zhao et al., 2024) None 105.1 - - -
Table 2: Comparing the quality and diversity of the generated video models. The generated videos consist of 25 frames (except for our nuScenes models which consist of 11 frames videos at 4 Hz) at a resolution of $312\times 520$, while the reported metrics from this table are evaluated at a resolution of $256\times 410$. The “# Cond. BBox” column reports the number of ground-truth input bounding box frames used by the generation pipelines. “None” indicates that no ground-truth frames are used, while “All” indicates that all ground-truth bounding box frames are utilized. If “# Cond. BBox” is $n\text{-to-}m$, then it represents the number of initial bounding box frames used by the pipeline is $n$ and the number of final bounding box frames used by the pipeline is $m$.
To assess the quality of video generation, we compare videos generated through 4 distinct pipelines:
  1. 1.
    Pre-trained Stable Video Diffusion (SVD) baselines [^1] Stable Video Diffusion (SVD) baseline is an image-to-video (I2V) model that generates a video sequence conditioned on a single video frame. without fine-tuning (initial frame $\rightarrow$ video)
  2. 2.
    Fine-tuned Stable Video Diffusion (SVD) baselines on the provided dataset (initial frame $\rightarrow$ video)
  3. 3.
    Teacher-forced Box2Video generation (initial frame and all bounding box frames $\rightarrow$ video)
  4. 4.
    bounding box generation with BBox Generator and Box2Video  (initial frame, one or three initial and one last bounding box frames $\rightarrow$ in-between bounding box frames and video).
We evaluate our generation across four metrics: Fréchet Video Distance (FVD) (Unterthiner et al., 2019), Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al., 2018), Structural Similarity Index Measure (SSIM) (Wang et al., 2004b) and Peak Signal-to-Noise Ratio (PSNR). These metrics either measure the consistency of frame pixels with the ground truth or the consistency of the frame latents extracted by another network. FVD [^2] FVD is highly sensitive to video configuration parameters—such as frame rate, clip duration, and spatial resolution—making direct comparisons of FVD values across studies challenging. Additionally, the metric’s sensitivity to sample sizes raises concerns, as some datasets may lack sufficient samples for convergence, leading to unreliable estimates. is an exception; it evaluates the generation distribution against the ground truth’s distribution. It is important to note that while many papers report their best-out-of-K results on these metrics, due to computational constraints, we evaluate our model on a single sample for each input.
The evaluated results are reported in Table 2 and visualizations are available in Appendix E.1. These results indicate that the generation quality improves as we condition on more ground-truth bounding box frames. Details regarding the metrics and their limitations are discussed in Appendix D.1.
@@ 4.4 Box2Video: Motion Control Evaluation
Figure 5: Illustrations of the Box2Video generations conditioned on ground truth 3D bounding box trajectories (2D for BDD) across various datasets. The 2D outlines of the ground-truth bounding boxes are overlaid on top.
@@
nuScenes Driving 25 0.661 0.833 0.734 0.381
Boximator [^3] Boximator only used the last generated frame’s bounding boxes to compute average percision scores. (Wang et al., 2024) MSR-VTT(Xu et al., 2016) Web videos 16 0.365 0.521 0.384 -
ActivityNet (Heilbron et al., 2015) Human-action 16 0.394 0.607 0.409 -
UCF-101 (Soomro et al., 2012) Human-action 16 0.212 0.343 0.205 -
@@
UCF-101 Soomro et al. (2012) Human-action 16 0.205 0.326 - -
Table 3: Average Precision scores obtained by comparing the YOLOv8 bounding box estimations of real and generated samples. Prior works (Wang et al., 2024; Li et al., 2024) do not report results on driving datasets; thus, we draw upon their reported performances on alternative datasets to provide a comparative context. The backbone model of Ctrl-V produces videos with 25 frames, while Boximator and TrackDiffusion create videos with 16 frames. Longer videos tend to have reduced quality and lower detection rates, which presents an extra challenge for our model (as it generates 56.25% more frames); yet it achieves greater precision compared to the other baseline models.
Our Box2Video is trained to control object motions through bounding boxes using a teacher-forcing approach, where only ground-truth bounding box frames are provided during the training phase. In this section, we analyze the fidelity of our Box2Video generations to the ground-truth bounding box conditions. To access the consistency of objects’ locations between our generated content and ground-truth, we compute the average precision of the bounding boxes in the generated frames and the ground-truth frames.
Average precision (AP) scores gauge the alignment of predicted/generated bounding boxes with the ground-truth labeling. In all related prior studies, average precision (AP) scores have been consistently reported. However, it is important to acknowledge that AP scores can vary across studies, depending on the specifics of the task setup. Boximator (Wang et al., 2024)’s motion control model predicts object locations in the scene, focusing solely on objects with consistent appearances across all frames. Their AP implementation disregards the object locations in the intermediate frames, comparing the objects’ locations only in the final frame. In contrast, TrackDiffusion (Li et al., 2024) uses TrackAP for evaluation, employing a QDTrack model (Fischer et al., 2023) to track instances in generated videos and comparing them to ground-truth labels. However, these evaluated datasets have limited instances, and TrackAP requires consistent tracking across frames, making it unsuitable for our project without modifications. Therefore, our AP score differs slightly from those in previous works.
First, we utilize the state-of-the-art object detection tool, YOLOv8 (Reis et al., 2024), to obtain the objects’ trackings from the generated and ground-truth scenes. Detailed information about the tool and our configurations is reported in Appendix D.3. Next, we match objects in each generated-vs-ground-truth frame pair based on spatial similarity – calculating the intersection over union (IoU) score to determine the similarity in location between objects’ bounding boxes. Our metric disregards object type and tracking IDs equivalence – assuming that objects close in location should naturally have the same type and IDs. Finally, we compute the average precision score following MS COCO protocol (Lin et al., 2015). Details are provided in Appendix D.4 and results are listed in Table 4.4. These results indicate that our Box2Video model is particularly adept at adhering to the specified conditions, especially when evaluated with a more lenient metric (i.e., a lower IoU threshold for the AP computation).